# On existence of neutral graph

### v1

@@
Abstract: Graph is considered neutral if its assortativity coefficient $r$ is equal to zero. In this paper, we address an outstanding conjecture, i.e., whether is there a neutral graph on $n$ vertices. First, we show that for $n\geq 7$, there is at least one neutral tree, which suggests that we find a representative of any order neutral graph. Additionally, we obtain that given $n\geq 13$, there exist at least one neutral non-tree graph.
@@ 1 Introduction
in which $d_{u}$ denotes degree of vertex $u$ and $e_{uv}$ represents an edge whose endpoints are vertices $u$ and $v$, separately. Graph $G$ is considered assortative if $r>0$, disassortative if $r<0$, and neutral otherwise. Obviously, it is easy to determine whether a given graph is neutral or not. However, to the best of our knowledge, there have not been published work to answer the following conjecture [3]-[6]
@@ 2.2 Main results
In the sequel, let us show main results composed of two subsections. The former shows existence of neutral tree, and the existence of neutral non-tree graph is displayed in the later.
@@ 2.2.1 Neutral tree
Lemma 2 Given graph $G$,
@@
Proof By means of Cauchy-Schwartz inequality, we prove Lemma 2. ∎
@@
Proof After some simple algebra, we complete Lemma 3. ∎
@@ 2.2.2 Neutral non-tree graph
Next, we will prove case $n\mod 6=2$. To this end, we need introduce a class of graphs. Without loss of generality, we use $G^{\oplus}$ to represent these graphs. Graph $G^{\oplus}$ is obtained based on two small graphs. Concretely speaking, given a neutral graph $G(1)=(V(1),E(1))$ and a $k$-regular graph $G(2)=(V(2),E(2))$, then we require that (i) $\beta$ incomplete edges [^2] An edge is considered incomplete if it just connects a vertex. be stemmed from each vertex of graph $G(2)$ and (ii) for each edge of graph $G(1)$, $(\alpha-1)$ incomplete edges be stemmed from every endvertex. In the meantime, we assume that $2(\alpha-1)|E(1)|=\beta|V(2)|$ holds true. Next, one merges an incomplete edge from graph $G(1)$ with another incomplete edge in graph $G(2)$ into an edge. Performing such manipulation until no incomplete edge is left. The resultant graph is $G^{\oplus}$.

### v2

# On the existence of neutral graph

@@
Abstract: Graph is considered neutral if its assortativity coefficient $r$ is equal to zero. In this paper, we address an outstanding conjecture, i.e., whether is there a neutral graph on $n$ vertices? First, we show that for $n\geq 7$, there is at least one neutral tree, which suggests that we find a representative of any order neutral graph. Additionally, we obtain that given $n\geq 13$, there exists at least one neutral non-tree graph.
@@ 1 Introduction
in which $d_{u}$ denotes degree of vertex $u$ and $e_{uv}$ represents an edge whose endpoints are vertices $u$ and $v$, separately. Graph $G$ is considered assortative if $r>0$, disassortative if $r<0$, and neutral otherwise. Obviously, it is easy to determine whether a given graph is neutral or not. However, to the best of our knowledge, there have not been published work answering the following conjecture [3]-[6]
@@ 2.2 Main results
In the sequel, let us show main results composed of two subsections. The former shows the existence of neutral tree, and the existence of neutral non-tree graph is displayed in the later.
@@ 2.2.1 Neutral tree
Lemma 2 Given an arbitrary graph $G$, the following expression holds
@@
Proof By means of Cauchy-Schwartz inequality, we can easily prove Lemma 2. ∎
@@
Proof Using some simple algebra, we complete Lemma 3. ∎
@@ 2.2.2 Neutral non-tree graph
Next, we will prove case $n\mod 6=2$. To this end, we need to introduce a class of graphs. Without loss of generality, we use $G^{\oplus}$ to represent these graphs. Graph $G^{\oplus}$ is obtained based on two small graphs. Concretely speaking, given a neutral graph $G(1)=(V(1),E(1))$ and a $k$-regular graph $G(2)=(V(2),E(2))$, then we require that (i) $\beta$ incomplete edges [^2] An edge is considered incomplete if it just connects a vertex. be stemmed from each vertex of graph $G(2)$ and (ii) for each edge of graph $G(1)$, $(\alpha-1)$ incomplete edges be stemmed from every endvertex. In the meantime, we assume that $2(\alpha-1)|E(1)|=\beta|V(2)|$ holds true. Next, one merges an incomplete edge from graph $G(1)$ with another incomplete edge in graph $G(2)$ into an edge. Performing such manipulation until no incomplete edge is left. The resultant graph is denoted by $G^{\oplus}$.